0 0
Add a comment Improve this question Transcribed image text
Answer #1

given energy E = 146 J

spring constant k = 106 N/m

so elastic potential energy = 1/2kx^2

1/2kx^2 = 146

x^2 = 2*146/106 = 2.75

x = sqrt(2.75) = 1.66 m

so option b is the correct answer